OpenStudy (anonymous):

Given That P=P=50/√X calculate Total revenue? and profit maximisation price and quantity.

OpenStudy (anonymous):

any one who can help me solve this?

OpenStudy (anonymous):

I really need help for this anyone?

OpenStudy (jlastino):

Can you rewrite the problem? I don't seem to figure out what needs to be done

OpenStudy (anonymous):

Given That P=50/√X calculate Total revenue? and profit maximisation price and quantity.Where P is the price.

OpenStudy (jlastino):

Are there given values for x?

OpenStudy (anonymous):

no.

OpenStudy (jlastino):

hmmm.. so you need to maximize the price...is this a calculus problem?

OpenStudy (anonymous):

yes

OpenStudy (jlastino):

Take the derivative (quotient rule) sqrt(x)(0)-(1/2(sqrtx))(50)/(sqrtx)^2 P' = -50/(2sqrt(x)*x)

OpenStudy (anonymous):

Jlastino can you help me solve this firm producing hokey sticks has a production function given by Q=2√(K.L) Where: K= Capital inputs L= Labour inputs In the short run, the firm amounts of capital equipments is Fixed at K=100. The rental rate for capital is V=1.00 Shillings and the wage rate for L=4.00 shillings. Calculate the firms short run costs functions as well as the short run average costs Functions. What is the firms short run marginal costs.